Attendees: Matt, Eric (Indiana), Elton, David, Elliott, Dan Carman, Blake, Rafael (JLab), Franz (CU), George, Kathryn, Zisis (Regina)
- Business Arising
- Action Items: See updated list below.
- Status Quo: The BCAL, carriage and cart are installed in the alcove, and the cart is bolted down to the floor in that location. All PMTs have been coupled and cabled, and the entire setup is covered with a layer of Tedlar. Kudos to Eric & IU for a phenomenal job on the cart, carriage and installation and to the Regina crew for the PMT work. Ditto for the JLab crew on electronics, cabling and DAQ. A H-B fileserver crash on Friday meant that HV control and collection of cosmic data is postponed until the server is repaired. This is expected later today or by tomorrow. In the meanwhile, David has found a different HV power supply so that work on testing the PMTs and looking for light leaks can continue.
- PMTs: It was agreed that caution should be applied in ramping up the HV for the PMTs until all light leaks have been sealed. The PMTs should be brought up gradually in, say, 100 V increments from 500 V and then stay at 1700 V until the cosmics checklist/protocol is completed. There is a one-channel NIM module in a RadPhi cabinet that can be used to expedite this effort. The issue of low energy deposited in cells versus individual and collective threshold was discussed. There does not seem to be a problem but we will be vigilant during pre-beam and beam conditions that may require electronics rewiring.
- Electronics and cabling: The fast core cable is in place and needs connectors. The CLAS folks have agreed to provide us with the required cables from the tagger, that will hoook up to a patch panel in the forward carriage. From then on those signals are our responsibility. Sergey Boyarinov and Sergey Pozdniakov will work on the H-B side of things.
- Trigger: David has placed information on the Trigger Supervisor in the HOWTO Wiki pages. There is also additional information on the Beam_Test_Trigger_and_electronics page.
- DAQ/online: Work will continue following the repair of the CLAS fileserver.
- Safety: We are awaiting signatures on the ESAD and TOSP to finalize this process. Eric will provide the missing text to Elliott for the outrigger displacement (close to cart).
- Run Plan: Zisis has been working on the Run Plan posted on the Wiki. He asked everyone to review it, contribute to it, or send comments. The key issues discusssed today were the overall goals, module positioning and priorities. Taking into account our goals, the BCAL+cart positioning constraints and general efficient operations, we are converging to the following priorities:
- Start with the BCAL on the platform position, with the beam impinging at its center and at normal incidence. Use the initial beam to check the performance of the BCAL and compare and calibrate it versus simulations. After the initial checks, complete the angular rotation scan at this point (~70-90o with respect to the beam).
- Move the BCAL to the alcove position. This is a ~2hr operation for the move, bolting plus interlock reset. Begin with a carriage setting that has the beam impinge at the maximum distance from one end of the BCAL that allows the full angle range scan (~10-45o). Repeat at closer points to the end (number of points to be specified in the future). Complete this objective by scanning the small angle range close to the end (~10-25o).
- Move the BCAL to the alcove position. This is a ~2hr operation for the move, bolting plus interlock reset. Measure at the two points closest to the ends of the BCAL (complete angular range in both, i.e. ~70-90o) and then continue at points on both sides, between the extremes and the middle, as time permits.
- Proposals from Athens to use a high-field-resistant Hamamatsu R3809U PMT in the place of one of the Burle ones and a pre-radiator in front of the BCAL will be discussed next week, after Zisis receives more info from Athens. Ditto for an idea from David to take measurements at different beam currents.
- Monte Carlo:
- Nothing new to report.
- Any other business
- The Wiki vs. e-log vs. hardbook log issue was discussed. It was agreed that all information that can be entered into the e-log should be, with the remainder logged in the hardbook in the counting room. Once the run begins, the Wiki sections related to the Beam Tests will be used mainly as a summary vehicle for activities and milestones as well as a document repository.
- BPM, Veto: Slow control cables and stand construction pending.
- Light Leak: Module has been painted and is covered with Tedlar. Light leak test pending in alcove.
- Tagger, Trigger, Cables: The fast core cable is pending. All other cabling is nearly complete.
- Data storage/Silo Request: David and Zisis will deal with this offline.
- Online Macros/MC Data steam/MC Validation & Binning: David and Zisis will work on these offline.
- Safety, COOP: TOSP and ESAD are nearing completion. COOP needs to be done ~2 weeks before the run.
- Cosmics Protocol: George and Zisis will circulate this soon.
- Run Plan: Decide on position/angle configurations and priorities.